OpenStudy (anonymous):

I need help with one more question, could you help me @TheSmartOne @ganeshie8?

RhondaSommer (rhondasommer):

i can help :)

RhondaSommer (rhondasommer):

so for this, i would recommend going to an online graohing calculator and graphing it out.

OpenStudy (solomonzelman):

2 by 4, versus 10 by 16

OpenStudy (solomonzelman):

If it is true that, 2/4 = 10/60 then they are "similar" (or "proportional")

OpenStudy (solomonzelman):

Oh excuse me

OpenStudy (solomonzelman):

10/16 (not 10/60

OpenStudy (solomonzelman):

Correction *** If it is true that, 2/4 = 10/16 then they are "similar" (or "proportional")

OpenStudy (solomonzelman):

but we can see that 16 is not twice as much as 10 (although 2/4 do result in 1/2), so ...

OpenStudy (solomonzelman):

Yes, therefore the two shapes are not similar/proportional.

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

Anytime!